# Shelf-Supervised Cross-Modal Pre-Training for 3D Object Detection

### v2

@@ 4 Experiments
In this section, we conduct extensive experiments to evaluate our proposed approach. We find that CM3D significantly improves over prior works in limited data settings (cf. Table 1). We include our code in the supplementary material, and will publicly release it to facilitate future research.
4.1 Experiment Setup
@@
4.2 Comparison to State-of-the-Art
We first evaluate the class-based zero-shot performance of CM3D to highlight the quality of our pseudo-labels. We benchmark class-agnostic performance in Appendix E. Next, we evaluate the performance of detectors trained on pseudo-labels. Lastly, we evaluate the performance of detectors pre-trained on CM3D psuedo-labels and fine-tuned on labeled data. Note that models are pre-trained with pseudo-labels from the entire training set and fine-tuned with ground truth labels from Kof the training set. We copy the results of prior work from CALICO and refer readers to [9] for further details. We present salient conclusions from Table 1 below.
Evaluating Zero-Shot Performance. We evaluate the quality of CM3D’s pseudo-labels (0% Training Data) and find that our method achieves 22.3% mAP and 22.1% NDS. We compare CM3D with SAM3D [61], a recently released zero-shot 3D detector. Notably, both CM3D and SAM3D use SAM [57] to group LIDAR points. However, CM3D groups LiDAR points on the image plane and SAM3D [61] groups LiDAR points in the 2D BEV plane. Our multi-modal zero-shot 3D detector significantly outperforms SAM3D’s LiDAR-only approach by 20.7%. Importantly, SAM3D was initially designed as a class-agnostic zero-shot detector. We posit that SAM3D’s poor performance can be attributed to nuScenes’ diverse classes. Unlike many prior methods that generate class-agnostic pseudo-labels [39, 40], our approach is able to generate multi-class predictions because we leverage images during training.
Distilling Multi-Modal Priors into Uni-Modal Models. Although CM3D requires paired RGB images and LiDAR sweeps for pseudo-label generation, we can easily train RGB-only or LiDAR-only student models (e.g. CenterPoint + CM3D). Importantly, these student models don’t require paired RGB-LiDAR data at inference (and are therefore compared against other models that only use LiDAR at inference). However, we find that the CenterPoint + CM3D (0% Training Data) student model performs worse than the pseudo-label generator (CM3D). This can be attributed to learning with noisy labels and not having access to the same sensor data as the teacher. Interestingly, we find that using multi-modal cues for both training and testing yields higher results.
Table 1: Semi-Supervised 3D Detection on NuScenes. In terms of zero-shot accuracy, pseudo-labels from CM3D outperforms prior art (SAM3D) by 20.7% NDS. With 5% supervised training data, pre-training CenterPoint with CM3D pseudo-labels improves over the prior art of PRC [9] by 8.1 mAP / 2.8 NDS. When comparing to Camera-LiDAR methods, pre-training BEVFusion with CM3D pseudo-labels outperforms prior art (CALICO) by 8.6 mAP / 4.6 NDS. We highlight the best LiDAR-only (L) results in blue, and the best Camera-LiDAR (L+C) results in red.
@@ 4.2 Comparison to State-of-the-Art
Significant Improvements in Low-Data Setting. Pre-training CenterPoint with CM3D pseudo-labels and fine-tuning on labeled data improves over the prior art of PRC [9] by 8.1 mAP and 2.8 NDS with 5% labeled data. Similarly, pre-training BEVFusion with CM3D pseudo-labels outperforms prior art (CALICO) by 8.6 mAP and 4.6 NDS. Notably all pre-training strategies improve over random initialization (Rand. Init.). This suggests that aligning your pre-training and fine-tuning tasks can yield better performance.
@@
Diminishing Returns with More Labeled Data. Although BEVFusion + CM3D shows significant performance improvements over CALICO in the low-data regime (e.g. 5% training data), the delta between both methods decreases as we increase the amount of training data (+8.6 mAP / +4.6 NDS at 5%, +3.3 mAP / +2.6 NDS at 10%, +1.4 mAP / +0.7 NDS at 20%, -0.3 mAP / -0.2 NDS at 50%). This suggests that the specific pre-training strategy matters less with increasing ground truth data.
Qualitative Results. We visualize the output of our pseudo-label generation approach in Fig. 4. Although our method generates reasonable predictions in many cases, we find that our method fails in cases of occlusions (where there is no 3D information) and in cases where the VLM predicts a false positive detection with high confidence. See Fig 4 and Appendix H for detailed analysis.
5 Conclusion
In this paper, we propose Cross-Modal 3D Detection Distillation (CM3D), a zero-shot approach for generating 3D bounding boxes using vision-language models. We demonstrate that pre-training detectors with zero-shot 3D bounding box pseudo-labels achieves higher downstream semi-supervised detection than prior contrastive learning objectives. We conduct comprehensive experiments to ablate our design choices and demonstrate that our simple method significantly improves over prior work in limited data settings. However, we find that our proposed pre-training strategy limits generalization to arbitrary downstream tasks. Moreover, our current pseudo-labeling pipeline requires LiDAR for precise box localization and HD maps for orientation estimation, which may not be readily available for broader robotics domains. We expand on the limitations and opportunities for future work in Appendix J.
6 Acknowledgements
We would like to thank Jenny Seidenschwarz for her feedback on early drafts. This work was supported in part by NSF GRFP (Grant No. DGE2140739).
Appendix A CM3D Pseudo-Label Refinement
Figure 5: Medoid Compensation. We find that all predicted medoids (shown in blue) tend to be radially biased toward the ego vehicle. This is because the LiDAR pointcloud only captures the visible surface of the car and not its full shape. To compensate for this bias, we “push” all predictions radially away, i.e., along the line connecting the center of the ego vehicle and the object medoid by a distance proportional to the object’s size. The corrected medoid is shown in yellow. Empirically, we show that this geometric trick improves mAP by 1.6% and NDS by 2.1%, respectively.
When training all detectors with pseudo-labels, we employ standard augmentation techniques (except for copy-paste augmentation, see Appendix D). Following established practices, we aggregate the past 10 sweeps for LiDAR densification using the provided ego-vehicle poses. We train CenterPoint and BEVFusion using the same hyperparameters prescribed by their respective papers. For CenterPoint, we first train the detector for 20 epochs with CM3D pseudo-labels and fine-tune the detector for 20 epochs using the limited set of annotations. For BEVFusion, we first pre-train the LiDAR-only branch using CM3D pseudo-labels for 20 epochs and fine-tune the LiDAR-only branch for 20 epochs using the limited set of annotations. We train the fusion model (RGB + LiDAR) for 6 epochs using the limited amount of labeled training data. Lastly, we fine-tune all models using self-training. Specifically, we use the fine-tuned model to generate new pseudo-labels on the unlabeled portion of the train-set and retrain the detector on the entire train-set (including the limited set of ground truth labels and pseudo-labels) from scratch. We ablate self-training further in Appendix B. We conduct all experiments on 8 RTX 3090 GPUs.
Ablation Study. We ablate our pseudo-label generation algorithm to determine how each component improves the baseline. As discussed above, we find that medoid compensation has the greatest impact on pseudo-label quality, improving mAP by 1.6% and NDS by 2.1%. Tuning the Detic text prompts by including synonyms also improves results significantly. Finally, the 3D distance-based NMS helps remove duplicates present in the overlapping regions of the ring cameras and increases the mAP by 1%.
Table 2: Ablation on Pseudo-Label Generation. We analyze the impact of each component over the baseline (cf. Fig 2). Importantly, we find that prompt engineering, medoid compensation, and non-maximum suppression have the greatest impact.
          Method           mAP $\uparrow$           NDS $\uparrow$
          Baseline           18.6           17.8
           + Prompt Engineering           19.7 (+1.1)           18.4 (+0.6)
           + LiDAR Accumulation           20.0 (+0.3)           18.6 (+0.2)
           + Mask Erosion           20.2 (+0.2)           19.1 (+0.5)
           + Medoid Compensation           21.8 (+1.6)           21.3 (+2.1)
           + Non-Maximal Suppression           22.8 (+1.0)           21.9 (+0.6)
           + Late Fusion           23.0 (+0.2)           22.1 (+0.2)
Appendix B Ablation on Self-Training
@@
Appendix C Ablation on Scaling Up Pseudo-Labels
Appendix D Ablation on Copy-Paste Augmentation
@@
Appendix E nuScenes Class Agnostic Evaluation
In the main paper, we present nuScenes detection results for CM3D pseudo-labels averaged over 10 classes. However, prior pseudo-label generation methods only evaluate class-agnostic performance [40]. For fair comparison with prior work, we evaluate CM3D pseudo-labels without differentiating between classes in Table 6. We find that our approach outperforms prior work by 14.5% AP and 0.6% NDS, highlighting the benefit of using foundational priors and multi-modal inputs. In addition, prior works use motion cues from scene flow to generate object proposals for moving objects [39]. In contrast, our approach localizes both static and moving objects and classifies them using VLMs.
Table 6: nuScenes Class Agnostic Evaluation. We evaluate class-agnostic performance of CM3D pseudo-labels for fair comparison with prior work. We find that our method significantly outperforms LISO-TF by 14.5% AP.
DBSCAN [25]   L   0.8   10.9   0.980   3.120   0.962
RSF [71]   L   1.9   18.3   0.774   1.003   0.507
Oyster-CP [72]   L   9.1   21.5   0.784   1.514   0.521
Oyster-TF [72]   L   9.3   23.3   0.708   1.564   0.448
LISO-CP [40]   L   10.9   22.4   0.750   1.062   0.409
LISO-TF [40]   L   13.4   27.0   0.628   0.938   0.408
CM3D (Ours)   L + C   27.9   27.6   0.592   0.872   0.428
@@ Appendix F Replacing Detic with GroundingDINO
Appendix G Comparing ChatGPT Priors vs. Real Shape Priors
@@ Appendix I Evaluating CM3D on Waymo
Table 12: Distilling Multi-Modal 3D Psuedo-Labels to RGB-based 3D Detectors. We train the camera-only branch of BEVFusion with CM3D pseudo-labels. Although distilling CM3D detections into BEVFusion-C performs considerably worse than BEVFusion-L + CM3D and BEVFusion + CM3D , we demonstrate that our shelf-supervised framework allows us to distill multi-modal information from expensive RGB + LiDAR sensors into models trained with cheaper RGB-only sensors.
  Training Data      Method Modality   mAP $\uparrow$   NDS $\uparrow$
  Train   Test
     CM3D   L + C   L + C   23.0   22.1
  0%      BEVFusion-L [62] + CM3D   L + C   L   16.7   21.4
  (Unsupervised)      BEVFusion-C [63] + CM3D   L + C   C   11.7   16.1
     BEVFusion [63] + CM3D   L + C   L + C   20.6   23.3
Future Work: Distilling Multi-Modal 3D Psuedo-Labels to RGB-based 3D Detectors. We repurposed our shelf-supervised framework to learn RGB-only 3D detectors, allowing us to distill multi-modal information from expensive RGB + LiDAR sensors into models trained with cheaper RGB-only sensors. Future work should explore ways of improving cross modal detection distillation of RGB + LiDAR into RGB-only models.
@@
Future Work: Application to Broader Robotics Tasks. Our approach leverages domain-specific insights, and may not be directly generalizable to robotics applications without LiDAR or HD Maps. Future work can adapt our pseudo-labeling framework for such situations. For example, instead of using LiDAR, one can directly use metric depth estimates [piccinelli2024unidepth] from state-of-the-art monocular depth estimators. Similarly, one can estimate orientation with rotating calipers or track-level information. We present preliminary results of our pseudo-labeler applied to non-AV applications in Fig. 9.
Figure 9: Generalized Pseudo-Labeler. We visualize the output of our generalized psuedo-labeler on several images above. Importantly, we don’t assume we have access to LiDAR for precise depth estimates or HD maps for orientation estimation. Instead, we use UniDepth [piccinelli2024unidepth] to generate a pseudo-point cloud and estimate orientation using rotating calipers. Importantly, orientation estimates can be up to 180 degrees off because this current setup cannot estimate heading. This suggests that our pseudo-labeler can generalize beyond autonomous vehicle settings with minor modifications. We expect that semi-supervised training can further improve the quality of predicted bounding boxes.

### v3

@@ 4 Experiments
In this section, we conduct extensive experiments to evaluate our proposed approach. We find that CM3D achieves state-of-the-art unsupervised detection performance and significantly improves over prior works in limited data settings (cf. Table 3). We copy the results of prior work from CALICO and refer readers to [9] for further details.
@@
Table 1: Pseudo-Label Refinement. We analyze the impact of each component over the baseline [10] (cf. Fig 3). Importantly, we find that prompt engineering, medoid compensation, and non-maximum suppression have the greatest impact. See Appendix B for additional details.
        Method         mAP $\uparrow$         NDS $\uparrow$
        Baseline (Re-implementation of [10])         18.6         17.8
         + Prompt Engineering         19.7 (+1.1)         18.4 (+0.6)
         + LiDAR Accumulation         20.0 (+0.3)         18.6 (+0.2)
         + Mask Erosion         20.2 (+0.2)         19.1 (+0.5)
         + Medoid Compensation         21.8 (+1.6)         21.3 (+2.1)
         + Non-Maximal Suppression         22.8 (+1.0)         21.9 (+0.6)
         + Late Fusion         23.0 (+0.2)         22.1 (+0.2)
Table 2: Class Agnostic Evaluation on nuScenes. We evaluate class-agnostic performance of CM3D pseudo-labels for fair comparison with prior work. We find that our method significantly outperforms LISO-TF by 14.5% AP.
DBSCAN [24]   L   0.8   10.9   0.980   3.120   0.962
RSF [66]   L   1.9   18.3   0.774   1.003   0.507
Oyster-CP [67]   L   9.1   21.5   0.784   1.514   0.521
Oyster-TF [67]   L   9.3   23.3   0.708   1.564   0.448
LISO-CP [39]   L   10.9   22.4   0.750   1.062   0.409
LISO-TF [39]   L   13.4   27.0   0.628   0.938   0.408
CM3D (Ours)   L + C   27.9   27.6   0.592   0.872   0.428
Table 3: Semi-Supervised 3D Detection on nuScenes. In terms of zero-shot accuracy, pseudo-labels from CM3D outperforms prior art (SAM3D) by 20.7% NDS. With 5% supervised training data, pre-training CenterPoint with CM3D pseudo-labels improves over the prior art of PRC [9] by 8.1 mAP / 2.8 NDS. When comparing to Camera-LiDAR methods, pre-training BEVFusion with CM3D pseudo-labels outperforms prior art (CALICO) by 8.6 mAP / 4.6 NDS. We highlight the best LiDAR-only (L) results in blue, the best Camera-only (C) results in green, and the best Camera-LiDAR (L+C) results in red.
@@
Pseudo-Label Refinement. We ablate our pseudo-label generation algorithm to determine how each component improves the baseline. As discussed above, we find that medoid compensation has the greatest impact on pseudo-label quality, improving mAP by 1.6% and NDS by 2.1%. Tuning the Detic text prompts by including synonyms also improves results significantly. Finally, the 3D distance-based NMS helps remove duplicates present in the overlapping regions of the ring cameras and increases the mAP by 1%.
Class Agnostic Evaluation. Prior pseudo-label generation methods only evaluate class-agnostic performance [39]. For fair comparison with prior work, we evaluate CM3D pseudo-labels without differentiating between classes in Table 2. Prior works use motion cues from scene flow to generate object proposals for moving objects [38]. In contrast, our approach localizes both static and moving objects and classifies them using VLMs. CM3D outperforms LISO-TF [39] by 14.5% AP and 0.6% NDS, highlighting the benefit of using foundational priors and multi-modal inputs.
Evaluating Zero-Shot Performance. We evaluate the quality of CM3D’s pseudo-labels (0% Training Data) and find that our method achieves 22.3% mAP and 22.1% NDS. We compare CM3D with SAM3D [60], a recently released zero-shot 3D detector. Notably, both CM3D and SAM3D use SAM [56] to group LIDAR points. However, CM3D groups LiDAR points on the image plane and SAM3D [60] groups LiDAR points in the 2D BEV plane. Our multi-modal zero-shot 3D detector significantly outperforms SAM3D’s LiDAR-only approach by 20.7%. Importantly, SAM3D was initially designed as a class-agnostic zero-shot detector. We posit that SAM3D’s poor performance can be attributed to nuScenes’ diverse classes.
Distilling Multi-Modal Priors into Uni-Modal Models. Although CM3D requires paired RGB images and LiDAR sweeps for pseudo-label generation, we can easily train an RGB-only or LiDAR-only student model (e.g., CenterPoint + CM3D). Importantly, these student models don’t require paired RGB-LiDAR data at inference (and are therefore compared against other models that only use LiDAR at inference). However, we find that the CenterPoint + CM3D (0% Training Data) student model performs worse than the pseudo-label generator (CM3D). This can be attributed to learning with noisy labels and not having access to the same sensor data as the teacher. Interestingly, we find that using multi-modal cues for both training and testing yields higher results.
Significant Improvements in Low-Data Setting. Pre-training CenterPoint with CM3D pseudo-labels and fine-tuning on labeled data improves over the prior art of PRC [9] by 8.1 mAP and 2.8 NDS with 5% labeled data. Similarly, pre-training BEVFusion with CM3D pseudo-labels outperforms prior art (CALICO) by 8.6 mAP and 4.6 NDS. Notably all pre-training strategies improve over random initialization (Rand. Init.). This suggests that aligning your pre-training and fine-tuning tasks can yield better performance. See Appendix C for additional results with 20% and 50% ground truth labels. Notably, we find that the impact of our proposed pre-training approach diminishes when training with additional ground truth data.
Qualitative Results. We visualize the output of our pseudo-label generator in Fig. 4. Although our method generates reasonable predictions in many cases, we find that our method fails in cases of occlusions (where there is no 3D information) and in cases where the VLM predicts a false positive detection with high confidence. See Fig 4 for detailed analysis and Appendix A and additional visuals.
5 Conclusion
In this paper, we propose Cross-Modal 3D Detection Distillation (CM3D), a zero-shot approach for generating 3D bounding boxes using vision-language models. We demonstrate that pre-training detectors with zero-shot 3D bounding box pseudo-labels achieves higher downstream semi-supervised detection than prior contrastive learning objectives. We conduct comprehensive experiments to ablate our design choices and demonstrate that our simple method significantly improves over prior work in limited data settings. However, we find that our proposed pre-training strategy limits generalization to arbitrary downstream tasks. Moreover, our current pseudo-labeling pipeline requires LiDAR for precise box localization and HD maps for orientation estimation, which may not be readily available for broader robotics domains (Appendix J).
6 Acknowledgements
We would like to thank Jenny Seidenschwarz for her feedback on early drafts. This work was supported in part by the NSF GRFP (Grant No. DGE2140739).
Appendix A More Qualitative Results
Appendix B CM3D Pseudo-Label Refinement
Figure 6: Medoid Compensation. We find that all predicted medoids (on the left) tend to be radially biased toward the ego vehicle. This is because the LiDAR pointcloud only captures the visible surface of the car and not its full shape. To compensate for this bias, we “push” all predictions radially away (on the right), i.e., along the line connecting the center of the ego vehicle and the object medoid by a distance proportional to the object’s size. Empirically, we show that this geometric trick improves mAP by 1.6% and NDS by 2.1%, respectively.
When training all detectors with pseudo-labels, we employ standard augmentation techniques (except for copy-paste augmentation, see Appendix E). Following established practices, we aggregate the past 10 sweeps for LiDAR densification using the provided ego-vehicle poses. We train CenterPoint and BEVFusion using the same hyperparameters prescribed by their respective papers. For CenterPoint, we first train the detector for 20 epochs with CM3D pseudo-labels and fine-tune the detector for 20 epochs using the limited set of annotations. For BEVFusion, we first pre-train the LiDAR-only branch using CM3D pseudo-labels for 20 epochs and fine-tune the LiDAR-only branch for 20 epochs using the limited set of annotations. We train the fusion model (RGB + LiDAR) for 6 epochs using the limited amount of labeled training data. Lastly, we fine-tune all models using self-training. Specifically, we use the fine-tuned model to generate new pseudo-labels on the unlabeled portion of the train-set and retrain the detector on the entire train-set (including the limited set of ground truth labels and pseudo-labels) from scratch. We ablate self-training further in Appendix D. We conduct all experiments on 8 RTX 3090 GPUs.
Appendix C Diminishing Returns with More Labeled Data
Table 4: Semi-Supervised 3D Detection on nuScenes. We evaluate the impact of CM3D pre-training with increasing annotation budget. We find that performance differences between methods (including random initialization) shrinks considerably. This suggests that the specific pre-training strategy matters less with increasing annotation budget. We highlight the best LiDAR-only (L) results in blue, and the best Camera-LiDAR (L+C) results in red.
  Training Data       Method Modality   mAP $\uparrow$   NDS $\uparrow$
  Train   Test
@@ Appendix C Diminishing Returns with More Labeled Data
Although BEVFusion + CM3D shows significant performance improvements over CALICO in the low-data regime (e.g. 5% training data), the delta between both methods decreases as we increase the amount of training data (+8.6 mAP / +4.6 NDS at 5%, +3.3 mAP / +2.6 NDS at 10%, +1.4 mAP / +0.7 NDS at 20%, -0.3 mAP / -0.2 NDS at 50%). This suggests that the specific pre-training strategy matters less with increasing ground truth data.
@@ Appendix D Ablation on Self-Training
Appendix E Ablation on Copy-Paste Augmentation
@@ Appendix F Replacing Detic with GroundingDINO
Appendix G Ablation on Scaling Up Pseudo-Labels
Appendix H Comparing ChatGPT Priors vs. Real Shape Priors
@@ Appendix I Evaluating CM3D on Waymo
Future Work: Application to Broader Robotics Tasks. Our approach leverages domain-specific insights, and may not be directly generalizable to robotics applications without LiDAR or HD Maps. Future work can adapt our pseudo-labeling framework for such situations. For example, instead of using LiDAR, one can directly use metric depth estimates [76] from state-of-the-art monocular depth estimators. Similarly, one can estimate orientation with rotating calipers or track-level information. We present preliminary results of our pseudo-labeler applied to non-AV applications in Fig. 9.
Figure 9: Generalized Pseudo-Labeler. We visualize the output of our generalized psuedo-labeler on several images above. Importantly, we don’t assume we have access to LiDAR for precise depth estimates or HD maps for orientation estimation. Instead, we use UniDepth [76] to generate a pseudo-point cloud and estimate orientation using rotating calipers. Importantly, orientation estimates can be up to 180 degrees off because this current setup cannot estimate heading. This suggests that our pseudo-labeler can generalize beyond autonomous vehicle settings with minor modifications. We expect that semi-supervised training can further improve the quality of predicted bounding boxes.